# Going beyond Landauer scattering theory to describe spatially-resolved non-local heating and cooling in quantum thermoelectrics

### v1

@@ Appendix D Boltzmann equation in general case of finite Coulomb repulsion with momentum-dependent relaxation
We can use the exact solution for $l(p)=l=\text{const}$ [30] to look in detail at the spatial dependence of phonon heating in this case. As $\delta f^{s}(p,x)$ in Eq. (63) has the same decay with distance $x$ for all $p$, we know that $\delta f^{s}(p,x)$ itself must decay monotonically with $x$, so ${\cal P}_{\text{ph}}^{\rm s}(x)$ alone cannot lead to heating or cooling spots. In contrast, outside the strong Coulomb limit, $e\varphi(x)$ exhibits decay on another length scale, the screening length. Thus the combined effect of ${\cal P}_{\text{ph}}^{\rm s}(x)$ and ${\cal P}_{\text{ph}}^{(\varphi)}(x)$ may still lead to heating or cooling spots.

### v2

@@ Appendix D Boltzmann equation in general case of finite Coulomb repulsion with momentum-dependent relaxation
We can use the exact solution for $l(p)=l=\text{const}$ [32] to look in detail at the spatial dependence of phonon heating in this case. As $\delta f^{s}(p,x)$ in Eq. (63) has the same decay with distance $x$ for all $p$, we know that $\delta f^{s}(p,x)$ itself must decay monotonically with $x$, so ${\cal P}_{\text{ph}}^{\rm s}(x)$ alone cannot lead to heating or cooling spots. In contrast, outside the strong Coulomb limit, $e\varphi(x)$ exhibits decay on another length scale, the screening length. Thus, the combined effect of ${\cal P}_{\text{ph}}^{\rm s}(x)$ and ${\cal P}_{\text{ph}}^{(\varphi)}(x)$ may still lead to heating or cooling spots.